# Improvement of Spiking Neural Network with Bit Planes and Color Models

### v2

@@ V-A Experimental settings
To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [25]. The model and experiment process were implemented using PyTorch [42] and SpikingJelly[43]. Color conversion algorithms are based on implementations from Kornia [44]. For the baseline method in the experiments, the spike rate coding is employed with a step size of 10. To facilitate optimization, we utilize surrogate gradient methods applied to all IF neurons using the arctan surrogate function $S(U)$ [25, 13] with the defined derivative $\partial S$ of:
$\frac{\partial S}{\partial U}=\frac{1}{\pi}\cdot\frac{1}{1+(\pi U)^{2}}$ (7)
@@
The optimization process is conducted with the Adam algorithm [45], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets. Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function:
$L(x,y)=\frac{\sum_{n=1}^{N}l_{n}}{N},$ (8)
$l_{n}=-\sum_{c=1}^{C}w_{c}\log\frac{\exp{x_{n,c}}}{\sum_{i=1}^{C}\exp{x_{n,i}}}y_{n,c}$
where $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, $C$ is the total amount of class where $i\in C$ and $c$ is the mini-classes.
@@ V-B Experimental Results
V-B1 The effect of bit plane coding on grayscale images
@@
Baseline Bit plane coding only Proposed method
MNIST[36] 98.62 97.87 98.93
KMNIST[46] 94.13 92.33 95.58
Fashion-MNIST[47] 88.83 86.66 90.19
TABLE III: Comparison of validation accuracy among three coding methods. Best value is highlighted in bold.
Dataset Analysis results of coding methods (bits)
Baseline Bit plane coding only Proposed method
AKC SE AKC SE AKC SE
MNIST[36] 721.5 0.53 772.8 0.56 744.4 0.55
KMNIST[46] 1093 0.70 1221 0.70 1150 0.70
Fashion-MNIST[47] 1384 0.86 1377 0.82 1381 0.84
TABLE IV: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE). Highest values of AKC are denoted in red and lowest are denoted in blue.
@@
First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information. For that, we employed three popular datasets of grayscale images, including MNIST[36], KMNIST[46], Fashion-MNIST[47]. The results of MNIST dataset were already shown in Table I. However, they are still included here to see the behaviors of the coding methods over different datasets. As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel. This process would produce eight bit-plane channels that represent the image information in binary form. In this part, we investigate three options, namely 1) the baseline model (spike coding only), 2) using bit plane coding only, 3) the proposed model (spike coding combined with bit plane coding).
@@
As demonstrated in Table III, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy. As mentioned, using only bit-plane coding resulted in a reduction in accuracy; however, the combination of spike coding with bit plane coding (i.e. the proposed method) provides gain in all datasets. For any method, the accuracy value is highest with the MNIST dataset and lowest with the Fashion-MNIST dataset. This could be because among these three datasets, MNIST is the simplest one while Fashion-MNIST is the most complex one. Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.4% (with KMNIST and Fashion-MNIST). That means the impact of bit plane coding is more prominent with more complex datasets.
Upon further analysis of spike distribution complexities for various coding methods, as presented in Table IV, several insights emerge.
First, we observe that the baseline method has the lowest complexity in MNIST and KMNIST datasets, while it has the highest complexity in the Fashion-MNIST dataset. We suspect that this difference stems from the nature of these datasets. Both MNIST and KMNIST primarily use binary pixel values, with zeros representing the background and ones indicating the presence of a character. This binary representation results in a simpler spike distribution. In contrast, Fashion-MNIST employs a wider range of pixel values to capture the textures of various clothing items, which introduces greater complexity into the spike distribution when using baseline coding methods. This observation suggests that more complex visual datasets may lead to even more pronounced differences in spike distribution when using these coding methods.
Additionally, by introducing a hybrid coding approach that combines baseline and bit plane coding methods, we hypothesize the potential existence of an optimal complexity threshold at which the SNN might achieve peak performance. This idea of an optimizable global maximum suggests that complexity thresholds could play a key role in determining SNN effectiveness. Based on our data, we propose that measuring spike distribution complexity could be crucial in tuning SNNs to perform optimally across different datasets and tasks.
V-B2 The effect of bit plane coding on color images
@@
Baseline Bit plane coding only Proposed method
CIFAR10[26] 70.69 37.28 73.49
CIFAR100[26] 38.57 11.28 42.15
Caltech101[48] 61.67 43.28 64.55
Caltech256[49] 30.04 21.75 41.51
EuroSAT[50] 84.39 79.87 88.48
Imagenette[51] 72.99 48.20 78.68
Food101[52] 19.62 6.29 36.95
GTSRB[53] 93.36 95.85 95.35
TABLE V: Average validation accuracy on various computer vision dataset of our proposed method.
Dataset Analysis results of coding methods (bits)
Baseline Bit plane coding only Proposed method
AKC SE AKC SE AKC SE
CIFAR10[26] 7530 0.995 6764 0.996 7190 0.995
CIFAR100[26] 7374 0.995 6710 0.996 7079 0.995
Caltech101[48] 310k 0.993 260k 0.992 279k 0.992
Caltech256[49] 307k 0.991 257k 0.990 276k 0.990
EuroSAT[50] 30.3k 0.953 22.9k 0.944 27.0k 0.949
Imagenette[51] 334k 0.987 282k 0.994 302k 0.992
Food101[52] 339k 0.989 288k 0.995 308k 0.993
GTSRB[53] 308k 0.902 2914 0.223 5791 0.019
TABLE VI: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE), values in thousands are represented with k (e.g., 22k = 22,000). Highest values of AKC are denoted in red and lowest are denoted in blue.
Dataset Analysis results of coding methods (bits)
AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE
CIFAR10[26] 7190 0.995 7086 0.992 6671 0.972 6634 0.927 6752 0.957 7083 0.986 6857 0.988
CIFAR100[26] 7079 0.995 6975 0.991 6606 0.974 6551 0.929 6660 0.955 6984 0.987 6769 0.985
Caltech101[48] 279k 0.992 269k 0.978 231k 0.957 241k 0.905 242k 0.929 268k 0.989 239k 0.971
Caltech256[49] 276k 0.990 265k 0.976 233k 0.957 243k 0.907 244k 0.933 244k 0.989 242k 0.972
EuroSAT[50] 27.0k 0.949 26.5k 0.968 25.2k 0.944 25.3k 0.993 25.8k 0.950 26.2k 0.932 25.9k 0.966
Imagenette[51] 302k 0.992 292k 0.989 267k 0.957 271k 0.901 276k 0.941 288k 0.979 277k 0.976
Food101[52] 308k 0.993 297k 0.990 273k 0.986 277k 0.908 280k 0.927 294k 0.975 281k 0.969
GTSRB[53] 5791 0.019 5706 0.006 17.6k 0.489 86.1k 0.462 92.9k 0.499 5704 0.005 5703 0.006
TABLE VII: Comparison of analysis results with our proposed coding method across all tested color models, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE) measured in bits, values in thousands are represented with k (e.g., 22k = 22,000). Highest values of AKC are denoted in red and lowest are denoted in blue.
@@
GTSRB[53] 95.35 51.36 70.68 79.18 80.86 51.78 51.78
Average 65.15 59.21 62.02 63.36 63.49 59.53 59.61
TABLE VIII: Comparison of validation accuracy among coding methods.
For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[26], CIFAR100[26], EuroSAT[50], Caltech101[48], Caltech256[49], Imagenette[51] (which is a subset of ImageNet[54]), GTSRB[53], and Food101[52]. The default color model of these datasets is RGB. Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels. So, among these datasets, CIFAR10[26], CIFAR100[26] and EuroSAT[50] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
Across all tested datasets, our proposed methods consistently outperform the baseline method, with the notable exception of the GTSRB[53] dataset, where bit-plane coding alone appears to yield the best performance without the need for combination with traditional spike coding. This observation suggests that while our methods generally enhance performance by improving the representation of spike distributions, certain datasets may favor different coding approaches depending on their visual complexity and specific characteristics. In the GTSRB[53] case, it seems that bit-plane coding alone is sufficient to capture the essential features for effective processing.
The analysis results presented in Table VI highlight the relationship between complexity and performance in different coding methods. Bit-plane coding consistently exhibits lower complexities across datasets, while traditional spike coding tends to produce higher complexity values. This outcome aligns with our hypothesis that complexity increases in visual datasets with greater detail and variation.
In further investigating the GTSRB[53] dataset, we observed a significant reduction in spike distribution complexity (over 99% when comparing baseline with proposed scheme and pure bit planes coding in Table VI) alongside improved model performance. Usually a decrease in complexity can sometimes indicate a loss of critical information, it may also suggest that the spike distribution has become more streamlined, capturing essential features more effectively. However, a complexity reduction at such large scale could also introduce elements of instability (shown in Table VII), as it can be challenging to discern whether the decreased complexity represents a beneficial simplification or a detrimental loss of information.
Consequently, while our findings indicate promising results, we suggest further exploration of complexity thresholds to ensure consistent performance across varying datasets.
V-B3 Impacts of color models
In this part, also we investigate two options, namely 1) the baseline method and 2) proposed method.
The accuracy results of different cases are shown in Table VIII. We can see that, when using RGB color model, the accuracy of the proposed method is always higher than that of the baseline. The gain ranges from about 2% (with GTSRB[53] dataset) up to 11% (with Caltech256[49] dataset), depending on the employed dataset. This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
@@ V-B3 Impacts of color models
Interestingly, the results show that RGB is not always the best color model. Among the eight datasets, RGB provides the best performance in three cases (namely CIFAR100[26], Caltech256[49], GTSRB[53]). CMY does not achieve any best results, whereas each of the other five color models achieves the best performance once. Especially, with Caltech101[48] dataset, the accuracy of HSV color model is 68.41%, which is much higher than that of RGB color model (64.55%).
When considering the average of accuracy values across all datasets, the RGB color model is significantly better than other color models (mentioned the last row of Table VIII) and even though there are performance degradation on GTSRB[53], averaged result across all dataset proven to at least out performed baseline method.
From the above discussion, the key findings of our research can be summarized as follows.
  • The proposed method provides benefits for both grayscale images and color images.
@@
  • The gain of the proposed method depends on the employed dataset. For practical color datasets, the gain is from 2% up to 11%.
@@
  • In terms of average accuracy value over all datasets, RGB is the best color model and CMY is the worst one.
  • However, considering individual datasets, RGB is not always the best color model. It provides the best performance in only three datasets.
  • Especially, with the Caltech101 dataset, the performance of HSV color model (68.41%) is much better than that of the RGB color model (64.55%).
VI Discussion
VI-A Complexity or entropy?
In this study, we employ two primary metrics, an approximated Kolmogorov complexity and Shannon entropy, to analyze the effects of bit-plane coding on the distribution of spike trains. These metrics were chosen because they are widely used in signal processing and information theory, making them well-suited for analyzing the coded spike and bit-plane signals. By treating both coded spikes and bit planes as types of signals, we believe that Kolmogorov complexity and Shannon entropy can provide valuable insights into the complexity of spike distributions. While we conducted limited analysis with Shannon entropy, our results indicate a positive correlation between Shannon entropy and Kolmogorov complexity. This suggests that Shannon entropy could serve as an alternative to Kolmogorov complexity for those interested in measurement and analysis. However, we find Kolmogorov complexity to be more effective, as it provides clearer distinctions between measurements when compared to Shannon entropy.
Despite achieving substantial insights through these metrics, we recognize that measuring complexity and entropy alone may not fully capture the intricacies of image data. Images inherently contain structural complexity and spatial information, which can be difficult to quantify and analyze. Focusing solely on spike signals as sequential data may overlook these aspects. A helpful analogy is the historical debate in physics regarding the nature of light, with some scientists viewing light as a wave and others as particles. It was only with the discovery of wave-particle duality through experiments like the double-slit experiment that a more comprehensive understanding emerged. Similarly, viewing spike signals solely as sequential signals could reveal certain patterns but may not capture the full complexity underlying the optimization process in image processing tasks. Currently, tools designed specifically for spike signal analysis that incorporate spatial information for image-based tasks are underexplored. This gap highlights the need for further research into analysis methods that can integrate both signal and spatial information for more complete insights.

### v3

@@ VI-A Experimental settings
To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [11]. The model and experiment process were implemented using PyTorch [45] and SpikingJelly[46]. Color conversion algorithms are based on implementations from Kornia [47]. For the baseline method in the experiments, the Poisson-distributed rate coding [48, 49] is employed with a step size $T=10$ where spike $s_{T,i}$ for i-th pixel at step $T$ with intensity $P_{T,i}$ can be interpreted as:
$s_{T,i}=\begin{cases}1&\text{if }random(0.0,1.0)\geq P_{T,i}\\ 0&\text{otherwise}\end{cases}$ (13)
To enable efficient training, surrogate gradient methods were applied to all IF neurons. We utilized the arctangent surrogate gradient function $S(U)$[11, 10]. Also, we adopted the standard cross-entropy loss function with a softmax activation, commonly adopted for classification tasks[45, 50, 51], computed using mean reduction across mini-batches.
@@
Inspired by traditional deep learning approach in model fine-tunning [50, 51], the optimization process is conducted with the Adam algorithm [52], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets.
@@ VI-B Experimental Results
VI-B1 The effect of bit plane coding on grayscale image datasets
TABLE IV: Comparison of validation accuracy among coding methods. Best performances on each dataset are denoted in bold. Performance gain when compared with the baseline is highlighted in green.
@@ VI-B1 The effect of bit plane coding on grayscale image datasets
Baseline Bit plane coding Proposed method
MNIST[42] 98.62 97.87 98.93 (+0.31)
KMNIST[53] 94.13 92.33 95.58 (+1.45)
Fashion-MNIST[54] 88.83 86.66 90.19 (+1.36)
Average 93.86 92.29 94.90 (+1.04)
@@
First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information. For that, we employed three popular datasets of grayscale images, including MNIST[42], KMNIST[53], Fashion-MNIST[54]. The results of MNIST dataset were already shown in Table III. However, they are still included here to see the behaviors of the coding methods over different datasets. As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel. This process would produce 8 bit-plane channels that represent the image information in binary form. In this part, we investigate three options,namely:
  • the baseline model (spike coding only),
  • using bit plane coding only,
  • the proposed model (spike coding combined with bit plane coding).
@@
As demonstrated in Table IV, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy. As mentioned, using only bit-plane coding resulted in a reduction in accuracy; however, the combination of spike coding with bit plane coding (i.e. the proposed method) provides gain in all datasets. For any method, the accuracy value is highest with the MNIST dataset and lowest with the Fashion-MNIST dataset. This could be because among these three datasets, MNIST is the simplest one while Fashion-MNIST is the most complex one. Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.45% (with KMNIST) and averaged at 1.04%. That means the impact of bit plane coding is more prominent with more complex datasets.
VI-B2 Comparison of SNN variants
TABLE V: Comparison of validation accuracy of the baseline and the proposed method on different SNN variants. Performance gain when compared with the baseline is shown in green.
Accuracy of coding methods per dataset(%)
MNIST Fashion-MNIST KMNIST CIFAR10 Average
SNN variant Base-line Proposed method Base-line Proposed method Base-line Proposed method Base-line Proposed method Base-line Proposed method
SEW-ResNet18 (ADD) 98.62 98.93 (+0.31) 94.13 95.58 (+1.45) 88.83 90.19 (+1.36) 70.69 73.49 (+2.80) 88.07 89.55 (+1.48)
SEW-ResNet18 (AND) 97.41 98.48 (+1.07) 87.14 89.14 (+2.00) 90.13 94.26 (+4.13) 62.58 70.95 (+8.37) 84.32 88.21 (+3.89)
SEW-ResNet18 (IAND) 98.39 98.76 (+0.38) 88.92 90.15 (+1.23) 93.23 94.90 (+1.67) 67.28 71.40 (+4.12) 86.96 88.80 (+1.84)
Spiking-ResNet18 98.25 98.74 (+0.49) 87.47 89.37 (+1.90) 93.23 95.06 (+1.83) 46.15 61.79 (+15.1) 81.28 86.24 (+4.96)
Spiking-ResNet34 69.02 92.77 (+23.8) 69.59 79.62 (+10.0) 37.45 72.23 (+34.8) 19.28 32.29 (+13.0) 48.84 69.23 (+20.4)
Spiking-ResNet50 13.45 46.22 (+32.8) 12.80 57.20 (+44.4) 10.44 23.87 (+13.4) 17.94 27.78 (+9.84) 13.66 38.77 (+25.1)
In addition, we also evaluated our spike coding scheme in different SNN variants. We employ only both grayscale and color image datasets in this part, including MNIST, KMNIST, Fashion-MNIST and CIFAR10. We benchmarked against multiple variants of Spiking ResNet[26] with LIF surrogate neurons, which have demonstrated competitive performance in prior work, along with AND and IAND variants of SEW-ResNet18[11].
In shown in Table V, the results reveal that in all of the tested variants, the proposed method consistently achieves a higher accuracy than the baseline. The performance gains range from 0.31% (in the case of SEW-ResNet18 (ADD) on MNIST dataset) to 44.40% (in the case of Spiking-ResNet50 on Fashion-MNIST). Also, Spiking-ResNet variants tend to underperform when compared with SEW-ResNet18 variants. The average accuracy values across all datasets show that SEW-ResNet18 (ADD) is the best variant. Fang et al. [11] also demonstrated that this variant usually achieves superior performance. Therefore, in subsequent experiments, we exclusively utilize this variant for benchmarking.
VI-B3 The effect of bit plane coding on color images
TABLE VI: Average validation accuracy on various computer vision dataset of our proposed method. Best performances are denoted in bold. Performance gain when compare with the baseline is highlighted in green.
@@ VI-B3 The effect of bit plane coding on color images
Baseline Bit plane coding Proposed method
CIFAR10[30] 70.69 37.28 73.49 (+2.80)
CIFAR100[30] 38.57 11.28 42.15 (+3.58)
Caltech101[55] 61.67 43.28 64.55 (+2.88)
Caltech256[56] 30.04 21.75 41.51 (+11.47)
EuroSAT[57] 84.39 79.87 88.48 (+4.09)
Imagenette[58] 72.99 48.20 78.68 (+5.69)
Food101[59] 19.62 6.29 36.95 (+17.33)
Average 53.99 35.42 60.83 (+6.84)
For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[30], CIFAR100[30], EuroSAT[57], Caltech101[55], Caltech256[56], Imagenette[58] (which is a subset of ImageNet[60]) and Food101[59]. The default color model of these datasets is RGB. Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels. So, among these datasets, CIFAR10[30], CIFAR100[30] and EuroSAT[57] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
The accuracy results of different cases are shown in Table VI. We can see that, the accuracy of the proposed method is always higher than that of the baseline. The gain ranges from the lowest of 2.80% (with CIFAR100) up to 17.33% (with Food101 dataset), depending on the employed dataset. This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
TABLE VII: Validation accuracy comparison of proposed method among color models.
@@
Average 60.83 60.43 60.79 61.09 61.01 60.64 60.74
Figure 4: Validation accuracy visualization of proposed method among color models.
VI-B4 Impacts of color models
In this part, we investigate the effect of color models on the performance of proposed method. The results are provided in Table VII and also visualized in Fig. 4. In this figure, the baseline model is also included for the sake of comparison. From Fig. 4, it can be seen that our proposed method using different color models is consistently better than the baseline. Also, in general, the variations in performances of different color models are small.
@@ VI-B4 Impacts of color models
From Table VII, the results show that RGB is not always the best color model. Among the eight datasets, RGB provides the best performance in only 2 cases (namely CIFAR100 and Caltech256). CMY does not achieve any best results, whereas each of the other five color models achieves the best performance once. Especially, with Caltech101 dataset, the accuracy of HSV color model is 68.41%, which is considerably higher than that of RGB color model (64.55%).
When averaging the accuracy values across all datasets, the HSL color model slightly outperforms the others for both coding method, as indicated in the last row of Table VII. The CMY color model, which does not achieve any best results, also has the lowest average performance. Further analysis reveals a distinctive characteristic of the HSL model that may contribute to its relatively stable performance. HSL encodes the hue component using a 0–359 degree representation, which results in a higher number of spike counts. Consequently, the maximum number of bits required for representation increases to $n_{bit}=9$. As theoretically formulated in Section V, this greater bit depth might facilitates a higher SNR ratio that enables better performance. In contrast, the CMY model operates within a more limited range of $[0,100]$ and thus needs only $n_{bit}=7$ bits to represent the color values. Based on the same logic of Section V, this smaller number of bits reduces the gradient magnitude along with signal term for model SNR and thus results in its lower performance.
To rigorously quantify the computational overhead introduced by our method with different color models, we measure the running times of an image batch on forward propagation for both baseline and the proposed method. As shown in Table VIII, the increases in running times depend on datasets and color models, ranging from 20.52% (with CIFAR-10 dataset and CMY color model) to 105.01% (with CIFAR-10 dataset and HSL color model). Averaging over all datasets, the increases are from 48.57% (with CMY color model) to 85.35% (with HSL color model).
Figure 5 shows the relationship between the average gain in accuracy and the average increase in processing time for different color models. We can see that any gain in performance requires some increase in processing time. Interestingly, while having the smallest performance gain, the CMY color model also has the lowest processing time increase. In contrast, the HSL color model has the greatest gain as well as the highest increase in processing time.
VI-C Discussion
Based on these observations, the key findings of our analysis are summarized as follows:
  • The proposed method yields performance improvements for both grayscale and color images by increasing gradient magnitude and improve model SNR during training process.
@@ VI-C Discussion
  • The gain of the proposed method depends on the employed dataset, ranging from 0.31% to 1.45% in small grayscale image dataset and 2.80% to 17.33% for practical colored image dataset with RGB-based variants.
  • Proposed method can be generalized for different SNN variants.
  • Though RGB is the default color model in existing datasets, it is not always the best one, providing highest results in only two datasets.
@@
  • Based on both theoretical foundation and in terms of average accuracy value over all datasets, HSL is the best color model and CMY is the worst one. Yet HSL provides the best performance in only one dataset.
TABLE VIII: Comparison of model training times between baseline and proposed method using different color models. The values are average training times per epoch.
Model Variants Processing time by dataset (ms)
CIFAR-10 CIFAR-100 Caltech101 Caltech256 EuroSAT Imagenette Food101 Average
Baseline 41.67 43.55 141.25 146.92 66.05 135.71 140.66 102.26
Proposed (RGB) 63.66 (+52.77%) 65.39 (+50.15%) 235.13 (+66.46%) 240.41 (+63.63%) 116.77 (+76.79%) 225.83 (+67.07%) 234.74 (+66.89%) 168.85 (+65.12%)
Proposed (CMY) 50.22 (+20.52%) 55.83 (+28.20%) 210.56 (+49.07%) 215.63 (+46.77%) 102.31 (+54.90%) 217.32 (+60.78%) 211.64 (+50.46%) 151.93 (+48.57%)
Proposed (YCbCr) 61.23 (+46.94%) 62.31 (+43.08%) 234.44 (+65.98%) 242.25 (+64.89%) 118.64 (+79.62%) 226.41 (+67.50%) 229.95 (+63.48%) 167.89 (+64.18%)
Proposed (HSL) 85.43 (+105.01%) 87.99 (+102.04%) 256.71 (+81.74%) 277.38 (+88.80%) 131.46 (+99.03%) 237.94 (+76.03%) 249.89 (+77.66%) 189.54 (+85.35%)
Proposed (HSV) 64.36 (+54.45%) 65.37 (+50.10%) 237.17 (+67.91%) 241.10 (+64.10%) 120.21 (+82.00%) 224.03 (+65.74%) 232.85 (+65.54%) 169.30 (+65.56%)
Proposed (CIE XYZ) 51.12 (+22.68%) 53.45 (+22.73%) 215.11 (+52.29%) 220.92 (+50.37%) 108.71 (+64.59%) 215.65 (+59.54%) 212.06 (+50.76%) 153.86 (+50.46%)
Proposed (CIE LAB) 50.95 (+22.27%) 52.96 (+21.61%) 213.86 (+51.41%) 219.80 (+49.61%) 105.93 (+60.38%) 217.89 (+61.20%) 215.71 (+53.36%) 153.87 (+50.47%)
Figure 5: Relationship between average gain in performance (higher is better) and average increase in processing time (lower is better) of different color models.
Our experimental results demonstrate that bit-plane coding can enhance the performance of SNNs. Despite the significant performance gains, the added computational complexity is a limitation of our method. As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications. Also, another limitation is that, similar to recent studies [26, 11], the experiments in our study are still based on software simulations. Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions. First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs, especially for different machine learning tasks. Second, it is obvious that the choice of color model may depend on the task in question, such as medical image analysis or traffic object recognition. So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks. In addition, the proposed method should be implemented on some hardware platforms and tested in practical scenarios.